FLUORESCENCE EXCITATION SPECTROSCOPY OF THE NH_{2} A^{2}A_{1}- X^{2}B_{1} SYSTEM IN THE REGION 290-430 NM

\underline{J. XIN}, Department of Physics and Engineering Technologies, Bloomsburg University, Bloomsburg, PA 17815; H. FAN, I. IONESCU, C. ANNESLEY, AND S. A. REID, Department of Chemistry, Marquette University, Milwaukee, WI 53233.

The fluorescence excitation spectrum of the NH_{2} ^{2}A_{1}-^{2}B_{1} system has been measured in the region from 290 to 430 nm under jet-cooled conditions using a pulsed discharge source. A total of fourteen bands involving the pure bending states (0,ν′_{2},0) were observed and analyzed, including ten bands not previously measured to our knowledge. The spectra are largely free from rotational perturbations at the low N values accessed in this experiment, and the molecular constants were obtained from a least squares fit to the determined spin-rovibronic term values. The derived constants are in good agreement with theoretical predictions incorporating the effects of orbital angular momentum on the spin-rotational fine structure.